From a group of 7 men and 6 women in a firm, five persons are to be selected to form a committee for a project so that at least 3 men are there on the committee. In how many ways can it be done?


1
Expert's answer
2020-07-06T19:39:28-0400

Solution:

Scenario 1: 3 men and 2 women


Number of ways to select 3 men: 7C3 = (7 x 6 x 5)/3! = (7 x 6 x 5)/(3 x 2 x 1) = 35

Number of ways to select 2 women: 6C2 = (6 x 5)/2! = (6 x 5)/(2 x 1) = 15

Thus, the number of ways to select 3 men and 2 women is 35 x 15 = 525.


Scenario 2: 4 men and 1 woman


Number of ways to select 4 men: 7C4 = (7 x 6 x 5 x 4)/4! = (7 x 6 x 5 x 4)/(4 x 3 x 2 x 1) = 35

Number of ways to select 1 woman: 6C1 = 6

Thus, the number of ways to select 4 men and 1 woman is 35 x 6 = 210.


Scenario 3: 5 men


Number of ways to select 5 men: 7C5 = (7 x 6 x 5 x 4 x 3)/5! = (7 x 6 x 5 x 4 x 3)/(5 x 4 x 3 x 2 x 1) = 42/2 = 21


Thus, the number of ways to select a 5-person committee with at least 3 men is:525 + 210 + 21 = 756.


answer : 756 ways .